# Embodying Control in Soft Multistable Grippers from morphofunctional co-design

### v1

@@ 2.3 Position and stiffness inverse design
Based on the model’s performance and its agreement with the FE simulations, we pose an inverse problem to co-design different DPF topologies capable of reaching a desired pre-programmed final position (embodied position control). Since the discrete states of the system are already determined by the stable states, the optimization algorithm only searches within the fully actuated configurations, making it computationally efficient. We utilized Bayesian optimization [49] to iterate over various numbers of segments, dome heights for each of the segments ($H_{i}$), unit separation ($\text{U}_{\text{sep}}$), and unit length ($\text{U}_{\text{L}}$) (see Fig. 2a for parameter reference). Two different objective functions are utilized: one based on position and targeting a specific tip coordinate (Position Design), and the second one targeting a specific tip coordinate while maximizing the finger’s stiffness (Position + Stiffness Design) (see section A.3).
@@
$\min_{H_{i},t,\text{Unit}_{L},\text{Unit}_{s}}\sqrt{\left(\text{Target}_{xy}-\text{Tip dis}\right)^{2}}$ (5)
@@
where the Tip dis represents the tip point shown in Fig. 2a (Tip), which coincides with the top node of the finger’s last unit. A domain constraint is imposed to guarantee that every dome unit is higher than the one behind ($H_{i+1}>H_{i}$), which in certain cases can lead to up to N possible stable states (N been the number of segments).
@@
Five different tip coordinate targets were tested to evaluate the performance of our model and the optimization algorithm. For each case, the coordinates in $x$ and $y$ were given to the algorithm (Target Position in Fig. 4a), and the optimization was run by incrementally adding one unit per run and calculating the objective function (Eq. 5). To reduce computational cost, a maximum of 8 units was defined (Fig. 4b), although more units can be added if the objective requires a longer finger geometry. The geometry with the minimum objective function value (Obj in Table S1) was selected and plotted to test if the desired position was achieved (see Fig. 4a). Geometric parameters, coordinate targets, and objective function values are reported in Table S1. The final stable state and its comparison with each target can be observed in Fig. S6. To validate our optimization algorithm, we 3D printed the five different geometries shown in Table S4 using NinjaFlex 85A and measured the final tip position after dome inversion. An error of less than 15% was measured for each sample, showing good agreement with the optimization algorithm and the lattice model.
@@
The topologies generated by this methodology exhibit deterministic behavior, meaning that kinematic states are consistently attained as long as the gripper’s geometric parameters remain unchanged (see Fig. 4c and Movie 2). Specifically, we cycle each gripper topology between the initial and fully activated states, verifying the gripper’s ability to achieve the same aperture and positioning across multiple cycles (Tests 1-3 in Fig. 4c). This characteristic enables robust robot design, as the bistable elements provide intrinsic kinematic control dictated by the gripper’s mechanical response. Notice that the pressure for activating the units is not controlled using sensors and closed-loop control, illustrating the embodied self-regulation from the designed multistability. Consequently, the position control inherent to the designed multistable structure serves as control set points that can be embodied in the robot’s geometry and achieved with a single, open-loop controlled actuator.
@@
Figure 4: DPF inverse design and tests a) Results for Target 1 and 3 predicted by the model after the optimization. b) Objective function vs number of segments. The minimum value is selected for every target position. c) The gripper attains different stable kinematical states that can be predicted with the proposed energy-based model and are transferred to 3D-printed samples. d) Restriction domain $R_{c}$ for stiffness optimization of DPF. e) Experimental comparison between position and position+stiffness optimization.
@@
Position + stiffness design: The finger’s stiffness is calculated by computing the work necessary to perturb it from a given (programmed) stable configuration. We determine the system’s total energy by solving:
$\Pi=\sum_{i=1}^{n}\left[E_{L}(q)+E_{NL}(q)+E_{T}(q)\right]-W(q),$ (6)
where the external work $W$ requires prior knowledge of the direction of the external force and its path. Given this limitation, we utilized an alternative approach by creating a circular domain of radius $R_{c}$, within which none of the nodes of our lattice model are permitted to enter. By imposing this domain constraint, we establish a method to determine the system’s state at a distance of $R_{c}$ from the fully activated stable state. To implement this alternative approach, we impose an optimization constraint using Lagrange multipliers [50], given by:
$\min_{q}\left(\sum_{i=1}^{n}E_{L}(q)+E_{NL}(q)+E_{T}(q)\right)-\lambda g(x(q))$ (7)
$g(x(q))=\max(0,R_{c}-||x(q)-C||)$
where $R_{c}$ is the radius and $C$ is the center of the circle, respectively. By determining the system’s energy as a function of $R_{c}$ (see Fig. 4d), we can maximize it to increase the system’s stiffness. Using this methodology, the geometry of the DPF can be optimized to target a specific position with maximum stiffness in the neighborhood of that stable state (see Fig. 4e). We can rewrite Eq. 5 as a weighted optimization given by:
$\min_{H_{i},\text{Unit}_{L},\text{Unit}_{s}}w_{1}\sqrt{\left(\text{Target}_{xy}-\text{Tip dis}\right)^{2}}+w_{2}\frac{1}{\bar{\Pi}\left(R\right)}$ (8)
@@
where $w_{1}$ and $w_{2}$ are the weights assigned to each objective, and $\bar{\Pi}(R)$ is the average energy for a range of $R_{c}$ values. The number of segments (N) is kept consistent with the results obtained from the previous optimization (Eq. 5), meaning that only $H_{i}$, $\text{Unit}_{L}$, and $\text{Unit}_{s}$ are considered to maximize stiffness. The results of the stiffness and position optimization are shown in Table S2. Moreover, we experimentally test Target 5 geometries produced by both optimization procedures (see Fig. S9b). By following this approach, we increased the stiffness of the Target 5 geometry by approximately 40% (see Fig. 4e), which consequently enhances the achievable grasping force.
@@ 2.4 Dome Phalanx Robot (DPR) performance
Figure 5: Dome Phalanx Robot (DPR) assembly and applications: Fingers 1 to 4 ($A_{1}$ $A_{4}$) can be changed for different applications and mechanical behaviors. a) Multistable robot with four different set points: (i) Stable state 1 with 0 inverted units, (ii) Stable state 2 with two inverted units, (iii) Stable state 3 with four inverted units, and (iv) Stable state 4 with six inverted units. b) $A_{1}$ and $A_{3}$ are optimized for position and stiffness, while $A_{2}$ and $A_{4}$ are optimized for position. $A_{2}$ and $A_{4}$ exert $F_{Dis}$, and $A_{1}$ and $A_{3}$ exert $F_{stt}$ ($F_{stt}>F_{Dis}$). c) Bistable + Metastable finger behavior: Strain energy and pressure vs. time response. d) Snapshots of embodied pick-and-place tasks: A single pressure input is given for the robot to pick the object (Commanded response) and then release it depending on the viscoelastic material response (Programmed response)(see Movie 5).
@@
To explore the capabilities of our design approach and the advantages of incorporating instabilities in soft robotics, we fabricated a soft robot (Dome Phalanx Robot - DPR) by combining four DPFs ($A_{1}$ to $A_{4}$), as shown in Fig. 5a. We assembled the four DPFs using 3D printing a rigid base that prevents interaction between the fingers, imposes appropriate boundary conditions considered by our model, and couples our robot to a UR3 robotic arm (see Movie 1). This robot can be assembled by either combining fingers with the same topology ($A_{1}=A_{2}=A_{3}=A_{4}$), resulting in uniform global functionality, or by using two sets of actuators ($A_{1}=A_{3}$ and $A_{2}=A_{4}$), which provides multiple functionalities within the same robot body.
@@
We demonstrate three applications using the proposed design methodology and FDM 3D printing to create multiple robot topologies. Specifically, we focus on three different applications: 1) A multistable robot with at least three set points/stable states (see Fig. 5a); 2) A configuration with two finger pairs that can exert different grasping forces but reach the same final position (see Fig. 5b); 3) A combination of bistable and metastable dome geometries (see Fig. 2c and d), where the viscoelastic response of the material allows us to encode pick and place robotic behavior into the robot geometry (see Fig. 5c and d).
@@
1) Multistable soft robot: We further leverage the characteristics of our multistable DPF by programming different set points within the same geometry. Using the optimization algorithm, we can program three stable states by minimizing for three discrete configurations (2, 4, and 6 active units). This design approach can be implemented by modifying the objective function as follows:
$\min_{H_{i},\text{Unit}_{L},\text{Unit}_{s}}w_{j}\sqrt{\left(\text{Target j}_{xy}-\text{Tip dis j}\right)^{2}}+w_{M+1}\frac{1}{\bar{\Pi}\left(R\right)}$ (9)
@@
where $M$ is the number of targeted stable states, and $j$ is the stable state number ($0<j\leq M$). By doing this, we can optimize for three target positions and maximum stiffness in the fully activated state (six inverted units). For this application, we impose the number of set points, the target tip position for each set point, and the number of units. While the model could be expanded to further modify the unit cell’s dimensions and determine the optimal number of units, for simplicity, we fixed the number of units at six to ensure at least two active domes per stable state. We tested our system by performing an inverse design with three different target positions (see section A.3.3 for reference), using the same methodology described previously (section 2.3). An average error of 12% was found between the desired positions and the three final stable states, as shown in Fig. 5a ((i) - (iv)).
@@
This methodology demonstrates the versatility of our DPF and provides an alternative approach to encoding multiple set points within the finger’s body by modulating its unit cell geometry. The local geometric modifications create more accessible stable states, all of which can be controlled using a single, open-loop driven pressure inlet per finger. Since the curvature and stiffness of the system are functions of the number of active units (see Fig. S9b), the grasping force and carrying load capacity of the DPR also become functions of the active domes. Consequently, the robot can also program different set points for force (see Movie 3).
@@
2) Multiple functional robot: For the second application, we utilized the results shown in section 2.3, where two different objective functions were considered to create two different finger topologies that achieve the same tip position but exert different grasping forces. Using the data from Table S1 and Table S2, we designed a robot arrangement where two fingers ($A_{1}$ and $A_{3}$) produce more grasping force than the others ($A_{2}$ and $A_{4}$). We tested the grasping force using a compliant mechanism composed of two bars that compress as the grasping force increases (see Fig. 5b). We characterized the mechanism’s force vs. aperture to measure the robot’s grasping force (see Movie 4). Our experiments show that the two different objective functions result in a force application of 1.18 times ($F_{Dis}$ and $F_{stt}$) while reaching the same position in their fully activated stable state.
@@
3) Embodied soft robotic tasks:For the third application, we incorporate the time variable into our system. By leveraging the viscoelastic response and geometry of our dome structures, we can pre-program the duration for which the structure remains in its activated state before returning to its original state. As demonstrated in section 2.1, we can modulate the geometry to create a fully metastable finger (see Fig. 2c). However, the same finger can exhibit both bistable and metastable behaviors by combining bistable and metastable dome unit cells. This approach allows the finger to be encoded with a specific set point and operate in the vicinity of that point by actuating the metastable unit cells [34]. By integrating these two behaviors, we can enable task planning, such as pick-and-place operations (see Movie 5), where the robot can pick an object by activating all units, move to a specific position, and release the object without additional interventions and with just one pressure input (see Fig. 5). We tested this application by designing a five-segment finger with three bistable and two metastable units. The set point with three active units is designed with $H_{1}=4$ mm, $H_{2}=4.5$ mm, and $H_{3}=5$ mm, while the metastable units are programmed to reset after approximately 5 seconds (see section A.3.4 for details). As shown in Fig. 5c, an object can be held while all units are inverted, even when the pressure is released from the system. However, after the metastable units reset, the object falls from the gripper. Using this approach, we can program a simple pick-and-place operation as demonstrated in Fig. 5d and Movie 5.
3 Discussion

### v2

@@ 2.3 Position and stiffness inverse design
Based on the model’s performance and its agreement with the FE simulations, we pose an inverse problem to co-design different DPF topologies capable of reaching a desired pre-programmed final position to achieve embodied position control. Since the stable positions of the system are already determined by the stable states, the optimization algorithm only searches within the fully actuated configurations, making it computationally efficient. We utilized Bayesian optimization [51] to iterate over various numbers of segments, dome heights for each of the segments ($H_{i}$), thickness of the dome ($t$), unit separation for each of the segments ($\text{U}^{i}_{\text{sep}}$), and unit length for each of the segments ($\text{U}^{i}_{\text{L}}$) (see Figure 2a for parameter reference). Two different objective functions are utilized: the first based on position and targeting a specific tip coordinate (Position Design), and the second one targeting a specific tip coordinate while maximizing the finger’s stiffness (Position + Stiffness Design; see section A.6).
@@
$\min_{H_{i},t,\text{U}^{i}_{L},\text{U}^{i}_{\text{sep}}}\left(\text{Target}_{xy}-\text{Tip}_{\text{dis}}\right)^{2}$ (6)
@@
$\hskip 14.22636pt\quad U^{i+1}_{\text{sep}}=U^{i}_{\text{sep}},\quad i=1,\dots,N$
$\hskip 14.22636pt\quad U^{i+1}_{\text{L}}=U^{i}_{\text{L}},\quad i=1,\dots,N$
@@
where the $\text{Tip}_{\text{dis}}$ represents the tip point (see Tip in Figure 2a), which coincides with the top node of the finger’s last unit. A domain constraint is imposed to guarantee that every dome unit is higher than the one in front ($H_{i+1}\leq H_{i}$), which in certain cases can lead to up to N possible stable states (N been the number of segments). Moreover, we impose constraints for all unit separations and unit lengths to be equal for simplicity and design domain to guarantee bistability on all units, based on the phase diagram shown in Figure S1b. However, different restrictions can be imposed depending on the DPF design purpose, and parameters can be individually tuned to achieve a large variety of positions and stable states (see Movie 3).
@@
Our algorithm was tested by defining five different tip coordinate targets using two materials with distinct properties (NinjaFlex 85A and Cheetah 95A) optimized to achieve the desired positions. For each case, the target coordinates in $x$ and $y$ were provided to the algorithm (Target Position in Figure 4a), and the optimization was performed by incrementally adding one unit per run and calculating the objective function (Equation 6). To reduce the computational cost, a maximum of 8 units was specified (Figure 4b), although additional units can be included if the objective requires a larger finger geometry. The geometry with the minimum objective function value (Obj in Tables S3 and S4) was selected and plotted to verify whether the desired position was achieved (see Figure 4a).
Geometric parameters, coordinate targets, and objective function values are reported in Table S3 and S4. The final stable state and its comparison with each target can be observed in Figure S11. To validate our optimization algorithm, we 3D printed the five different geometries shown in Table S9 using both tested materials (NinjaFlex 85A and Cheetah 95A) and measured the final tip position after commanded dome inversion. An average error of 7.3% was measured across all samples, showing good agreement with the optimization algorithm and the lattice model.
@@
The topologies generated by this methodology exhibit deterministic behavior despite the use of polymeric materials, meaning that kinematic states are consistently attained as long as the gripper’s geometric parameters remain unchanged (see Figure 4c and Movie 4). Specifically, we cycle each gripper topology between the initial and fully activated states, verifying the gripper’s ability to achieve the same aperture and positioning across multiple cycles (Tests 1-3 in Figure 4c). Furthermore, we assessed material and time degradation effects by performing 300 cycles over three weeks (see Figure S17), which demonstrates that even after numerous activations and an extended period, the gripper’s performance remains within 5% of the initial design behavior. This characteristic enables robust robot design, as the bistable elements provide intrinsic kinematic control dictated by the gripper’s mechanical response. Notice that the pressure for activating the units is not controlled using sensors and closed-loop control, illustrating the embodied self-regulation from the designed multistability. Consequently, the position control inherent to the designed multistable structure serves as control set points that can be embodied in the robot’s morphology and achieved with a single, open-loop controlled actuator.
@@
Figure 4: DPF inverse design and experimental validation. (a) Predicted results for Target 1 and Target 3 from the optimization model (Position Design). (b) Objective function values as a function of the number of segments, with the minimum value identified for each target position. (c) The gripper achieves multiple stable kinematical states predicted by the energy-based model, which are successfully transferred to 3D-printed prototypes. (d) Stiffness determination of the DPF using a follower force approach, with Normalized strain energy vs Displacement curves shown for Objectives 1, 3, and 5. (e) Comparison of the proposed model and experimental results for position-only (Pos) optimization and combined position + stiffness optimization (Stf).
@@
Position + stiffness design: The finger’s stiffness is calculated by using the dynamic model described in section A.4, by perturbing the structure from a programmed stable configuration (i.e., set point). We determine the system’s total internal force by solving the dynamic equation:
$[M]\ddot{x_{i}}+F_{d}^{int}(x_{ij},\dot{x_{i}},\dot{x_{j}})+F_{d}^{iso}(\dot{x_{i}},\dot{x_{j}})+F_{\text{in}}=F_{ext}(x_{i,j},t)$ (7)
$F_{\text{in}}=\nabla\left(E_{L}(x_{ij})+E_{NL}(x_{ij})+E_{T}(x_{ij})\right)$
where $F_{\text{ext}}(x_{i,j},t)$ is a follower force applied at the tip of the DPF (see Figure 4) to perturb the DPF from its equilibrium position. The gradient of the internal forces in the structure, $F_{\text{in}}$, is obtained through symbolic differentiation of the expression described in Section 2.2. The force is applied at a constant load rate of 5 mm/min, ensuring a quasi-static regime. The stiffness of the DPF in the desired state is calculated by determining the internal force ($F_{\text{in}}$) and the corresponding tip displacement over time (see Figure 4). A linear fit is then applied to the resulting data. Using this methodology, the DPF’s geometry can be optimized to achieve a specific target position with maximum stiffness in the vicinity of that stable state (see Figure 4e). Equation 6 can thus be reformulated as a weighted optimization problem expressed as:
$\min_{H_{i},t,\text{Unit}^{i}_{L},\text{Unit}^{i}_{s}}w_{1}\left(\text{Target}_{xy}-\text{Tip}_{\text{ dis}}\right)^{2}+w_{2}\left(\frac{K_{\text{dis}}}{dF_{\text{in}}/dx}\right)^{2}$ (8)
@@
$\hskip 14.22636pt\quad U^{i+1}_{\text{sep}}=U^{i}_{\text{sep}},\quad i=1,\dots,N$
$\hskip 14.22636pt\quad U^{i+1}_{\text{L}}=U^{i}_{\text{L}},\quad i=1,\dots,N$
where $w_{1}$ and $w_{2}$ are the weights assigned to each objective, and $K_{\text{dis}}$ is the stiffness obtained from position inverse design (Equation 6). The number of segments (N) is kept consistent with the results obtained from the previous optimization (Equation 6), meaning that $H_{i}$, $\text{Unit}^{i}_{L}$, $t_{\text{lim}}$ and $\text{Unit}^{i}_{s}$ are considered to maximize stiffness. The results of the stiffness and position optimization are shown in Table S6 for Cheetah 95A and in Table S5 for Ninjaflex 85A. Moreover, we experimentally test Target 5 geometries produced by both optimization procedures (see Figure S16 and Movie 5). By following this approach, we had increased the DPF initial stiffness on average 1.94 times for all cases, achieving maximum stiffnesses of 0.55 N/mm, 0.3 N/mm, and 0.15 N/mm for cases 2, 4, and 5 respectably (see Figure 4e for case 5), which consequently enhances the achievable grasping force.
@@ 2.4 Dome Phalanx Robot (DPR) performance
To explore the capabilities of our design approach and the advantages of leveraging multistability in soft robotics, we fabricated a soft robot (Dome Phalanx Robot - DPR) by combining four DPFs ($A_{1}$ to $A_{4}$), as shown in Figure 5a. We assembled the four DPFs using 3D printing a rigid base that prevents interaction between the fingers, imposes appropriate boundary conditions our model considers, and couples our robot to a UR3 robotic arm (see Movie 6). This robot can be assembled by either combining fingers with the same topology ($A_{1}=A_{2}=A_{3}=A_{4}$), resulting in uniform global functionality, or by using two sets of actuators ($A_{1}=A_{3}$ and $A_{2}=A_{4}$), which provides multiple functionalities within the same robot body.
@@
We demonstrate different operation modes using the proposed design methodology and FDM 3D printing to create multiple robot topologies. Specifically, we focus on three operation modes: 1) A multistable robot with at least three set points/stable states (see Figure 5a); 2) A configuration with two finger pairs that can exert different grasping forces but reach the same final position (see Figure 5b); and 3) A combination of bistable and metastable dome geometries (see Figure 2c and d), where the DPR’s constitutive material viscoelastic response of the material allows us to embody a pick and place robotic behavior into the robot geometry (see Figure 5c and d and Movie 7). For each of these operation modes, different constraints (detailed below) are imposed on the model to ensure that the structure is accurately designed for each specific task.
Figure 5: Dome Phalanx Robot (DPR) assembly and modes: Fingers 1 to 4 ($A_{1}$ $A_{4}$) can be changed for different operation modes and mechanical behaviors. a) Multistable robot with four different set points: (i) Stable state 1 with 0 inverted units, (ii) Stable state 2 with two inverted units, (iii) Stable state 3 with four inverted units, and (iv) Stable state 4 with six inverted units. b) $A_{1}$ and $A_{3}$ are optimized for position and stiffness, while $A_{2}$ and $A_{4}$ are optimized for position. $A_{2}$ and $A_{4}$ exert $F_{Dis}$, and $A_{1}$ and $A_{3}$ exert $F_{\text{stt}}$ ($F_{\text{stt}}>F_{\text{Dis}}$). c) Bistable + Metastable finger behavior: Strain energy and pressure vs. time response. d) Snapshots of embodied pick-and-place tasks: A single pressure input is given for the robot to pick the object (Commanded response) and then release it depending on the viscoelastic material response (Programmed response)(see Movie 7).
@@
1) Multistable soft robot: We further leverage the characteristics of our multistable DPF by programming different set points for the same robot morphology (i.e., DPFs geometries). Using the optimization algorithm, we can program three stable states by minimizing for three discrete configurations (2, 4, and 6 active units). This design approach can be implemented by modifying the objective function as follows:
$\min_{H_{i},t,\text{Unit}^{i}_{L},\text{Unit}^{i}_{s}}w_{j}\left(\text{Target j}_{xy}-\text{Tip}_{\text{dis j}}\right)^{2}+w_{M+1}\left(\frac{K_{\text{dis}}}{dF_{\text{in}}/dx}\right)^{2}$ (9)
@@
$\hskip 14.22636pt\quad U^{i+1}_{\text{sep}}=U^{i}_{\text{sep}},\quad i=1,\dots,N$
$\hskip 14.22636pt\quad U^{i+1}_{\text{L}}=U^{i}_{\text{L}},\quad i=1,\dots,N$
@@
where $M$ is the number of targeted stable states, and $j$ is the stable state number ($0<j\leq M$). By doing this, we can optimize for three target positions and maximum stiffness in the fully activated state (six inverted units). For this design, we impose the number of set points, the target tip position for each set point, and the number of units. It is worth noting that for this case, there is a different constraint for the dome height (i.e., $H_{i+1}>H_{i}$), which allows each of the states to be useful for grasping objects while being accessible in a sequential manner. While the model could be expanded to modify the unit cell’s dimensions further and determine the optimal number of units, for simplicity, we fixed the number of units at six to ensure at least two active domes per stable state. We tested our system by performing an inverse design with three different target positions (see section A.6.3 for reference), using the same methodology described previously (section 2.3). An average error of 12% was found between the desired positions and the three final stable states, as shown in Figure 5a ((i) - (iv)).
@@
This methodology demonstrates the versatility of our DPF and provides an alternative approach to encoding multiple set points within the finger’s body by modulating its unit cell geometry. The local geometric modifications create more accessible stable states, which can be controlled using a single, open-loop driven pressure inlet per DPF. Since the systems’s curvature and stiffness are functions of the number of active units (see Figure S15b), the grasping force and carrying load capacity of the DPR also become functions of the active domes. Consequently, the robot can also be programmed with different set points for force (see Movie 6). Moreover, to showcase the gripper’s versatility to grasp various objects and its carrying capacity, we performed grasping of different objects on different stable states of the DFR (see Figure S18 and Movie 6). Notably, we observed that the restriction of the object can reset some of the dome units, which makes the DFP adapt to the object shape and perform a better grip depending on the object size. A maximum load-to-weight ratio of 10.25 is determined by loading the robot until it can not withstand the weight (see Movie 6).
@@
2) Force-adaptable robot: For the second operation mode, we utilized the results shown in section 2.3, where two different objective functions were considered to create two different finger topologies that achieve the same tip position but exert different grasping forces. Using the data from Table S3 and Table S5, we designed a robot arrangement where two fingers ($A_{1}$ and $A_{3}$) produce more grasping force than the others ($A_{2}$ and $A_{4}$). We tested the grasping force using a compliant mechanism composed of two bars that compress as the grasping force increases (see Figure 5b). We characterized the mechanism’s force vs. aperture to measure the robot’s grasping force (see Movie 5). Our experiments show that the two different objective functions result in a force increase of 1.8 times ($F_{{\text{Dis}}}$ and $F_{{\text{stt}}}$) while reaching the same position in their fully activated stable state.
@@
3) Embodied soft robotic tasks: In the third operation mode, we leverage our ability to program our system’s dynamics. By leveraging our dome structures’ viscoelastic response and geometry, we can pre-program the duration for which the structure remains activated before returning to its original state. As demonstrated in section 2.1, we can modulate the geometry to create a fully metastable finger (see Figure 2d and e). However, the same finger can exhibit bistable and metastable behaviors by combining bistable and metastable dome units. This approach allows for encoding a desired set of points into the finger and operate in the vicinity of that point by actuating the metastable unit cells (See Movie 2). Combining the meta- and bi-stability of the units enables us to design task planning operations (such as pick-and-place; see Movie 7), where the robot can pick an object by activating all units, move to a specific position, and release the object without additional interventions and with just one pressure input (i.e., eliminating the need for a release command; see Figure 5d). We tested this operation mode by designing a six-segment finger with four bistable and two metastable units. The set point with four active units is designed with $H_{1}=4$ mm, $H_{2}=4.5$ mm,$H_{3}=5$, and $H_{4}=5$ mm, while the metastable units are designed to maximize the resetting time for two different loading times (section A.6.4). We approximate the loading profile, by measuring the pressure experimentally and approximating it as shown in Figure S13. The obtained results show how an object can be held while all units are inverted, even when the system releases pressure (Figure 5c). However, the object is automatically released from the gripper after the metastable units reset, as desired from the programmed dynamics. Using this approach, we can program a simple pick-and-place operation by activating all domes or perform a simple pinch operation as demonstrated in Figure 5d and Movie 7. Further flexibility in the attainable behavior of our robots can be obtained by including the actuation loading time, which results in the same DPR morphology adapting its response to can pick an object and release it at two different times as demonstrated in Movie 7.
3 Conclusion